OpenStudy (anonymous):

Simplify the given expression to radical form, justify each step by identifying the properties of rational exponents used. All work must be shown. x to the two–thirds power, over x to the four–ninths power

OpenStudy (tkhunny):

\(\dfrac{x^{2/3}}{x^{3/9}}\) Do you have this rule? \(\dfrac{a^{b}}{a^{c}} = a^{b-c}\)?

OpenStudy (anonymous):

how

OpenStudy (tkhunny):

Just like it is shown. \(\dfrac{3^{5}}{3^{2}} = 3^{5-2} = 3^3\)